A horse drawn carriage travels 47(2/3) kilometers in 4(1/3) hours. At this rate how many kilometers does it travel per hour
Answer by stanbon(75887)

A horse drawn carriage travels 47(2/3) kilometers in 4(1/3) hours. At this rate how many kilometers does it travel per hour
------
rate = distance/time
-----
rate = (47 2/3 km)/(4 1/3 hrs) = (143/3)/(13/3) = 143/13 = 11 km/hr
--------------------
Cheers,
Stan H.
---------------